OpenStudy (anonymous):

help please There are 8 people on the ballot for regional judges. Voters can vote for any 4. Voters can choose to vote for 0¸ 1¸ 2¸ 3¸ or 4 judges. In how many different ways can a person vote? A. 5 B. 56 C. 70 D. 163

OpenStudy (anonymous):

answer = 8C0 + 8C1 + 8C2 + 8C3 + 8C4 = 1 + 8 + 28 + 56 + 70 = 163

OpenStudy (anonymous):

thank you :)